# Speak & Improve Corpus 2025: an L2 English Speech Corpus for Language Assessment and Feedback

### v1

@@ Abstract
We introduce the Speak & Improve Corpus 2025, a dataset of L2 learner English data with holistic scores and language error annotation, collected from open (spontaneous) speaking tests on the Speak & Improve learning platform. The aim of the corpus release is to address a major challenge to developing L2 spoken language processing systems, the lack of publicly available data with high-quality annotations. It is being made available for non-commercial use on the ELiT website. In designing this corpus we have sought to make it cover a wide-range of speaker attributes, from their L1 to their speaking ability, as well as providing manual annotations. This enables a range of language-learning tasks to be examined, such as assessing speaking proficiency or providing feedback on grammatical errors in a learner’s speech. Additionally the data supports research into the underlying technology required for these tasks including automatic speech recognition (ASR) of low resource L2 learner English, disfluency detection or spoken grammatical error correction (GEC). The corpus consists of around 340 hours of L2 English learners audio with holistic scores, and a subset of audio annotated with transcriptions and error labels.
@@ 1 Introduction
Spoken language assessment and feedback are crucial components of language learning. Developing robust and inclusive automated tools for these tasks remains a significant challenge. One important limitation to developing systems is the lack of high-quality labelled data. To address this issue and encourage research in these areas, we are distributing a new annotated corpus of spoken L2 (second language) learner English, collected from speaking tests performed on the Speak & Improve (S&I) learning platform [1]. The Speak & Improve Corpus 2025 consists of around 340 hours of recordings of L2 English learner speech on open speaking tasks, annotated with English speaking proficiency scores. In addition a subset of the data, around 60 hours, was manually transcribed, including disfluencies, and grammatically error corrected phrases [2]. The S&I platform was used by speakers from around the globe yielding a wide range of L1 (first language) backgrounds. The proficiency levels of the speakers range from Elementary (A2) to Advanced (C1) on the CEFR scale [3]. This breadth of speaker attributes and abilities is crucial for the development of inclusive language learning technologies. We believe this is the most comprehensive L2 English learner public corpus released to date.
@@ 4.2 Data distribution and selection
The Dev and Eval test sets were first defined to consist of full test submissions only i.e. where the learner has completed the test and all parts were successfully annotated at each phase of interest. 300 fully annotated test submissions were selected for each of the Dev and Eval test sets. These were extended for spoken language assessment (SLA) with a further (approximately) 150 submissions which had undergone phase 1 holistic proficiency scoring only. The submissions were selected so that the grade distribution was as even as possible but also reflects the nature of the learners using the platform. Table 4 shows the grade distribution for the Dev set. As can be seen the majority are in the CEFR range B1-B2+. The Eval and Train sets have similar grade distributions. There was insufficient L1 and speaker data provided to incorporate this information into the data set selection. Note, this does mean that a speaker could appear in both the training and test sets. Table 5 shows the statistics for the Dev and Eval sets.
@@
Submissions Utterances Trans GEC All Transcript   GEC
Dev 438 5616 24.7 22.5 38.5 140k 105k
Eval 442 5642 24.5 22.0 38.7 140k 104k
@@
Submissions Utterances Trans GEC All Transcript   GEC
1 3068 18072 7.7 4.2 83.6 47k 22k
3 3060 3060 6.0 3.1 51.0 37k 14k
4 3005 3005 6.0 1.6 50.1 36k 7k
5 3085 15353 9.7 4.9 85.3 60k 24k
Total 6640 39490 28.4 12.8 270.0 170k 65k
@@ 5 Conclusions and Future Work
The S&I Corpus 2025 consists of audio and manual annotations of open speaking test submissions performed on the Speak & Improve learning platform. We provide around 340 hours of speech data from L2 English speakers at A2 to C1 CEFR levels, split into train, dev and eval data sets. There are around 950 fully annotated test submissions and the equivalent of a further 2500+ test submissions with manual speaking proficiency assessment scores.

### v2

@@ Abstract
We introduce the Speak & Improve Corpus 2025, a dataset of L2 learner English data with holistic scores and language error annotation, collected from open (spontaneous) speaking tests on the Speak & Improve learning platform. The aim of the corpus release is to address a major challenge to developing L2 spoken language processing systems, the lack of publicly available data with high-quality annotations. It is being made available for non-commercial use on the ELiT website. In designing this corpus we have sought to make it cover a wide-range of speaker attributes, from their L1 to their speaking ability, as well as providing manual annotations. This enables a range of language-learning tasks to be examined, such as assessing speaking proficiency or providing feedback on grammatical errors in a learner’s speech. Additionally the data supports research into the underlying technology required for these tasks including automatic speech recognition (ASR) of low resource L2 learner English, disfluency detection or spoken grammatical error correction (GEC). The corpus consists of around 315 hours of L2 English learners audio with holistic scores, and a subset of audio annotated with transcriptions and error labels.
@@ 1 Introduction
Spoken language assessment and feedback are crucial components of language learning. Developing robust and inclusive automated tools for these tasks remains a significant challenge. One important limitation to developing systems is the lack of high-quality labelled data. To address this issue and encourage research in these areas, we are distributing a new annotated corpus of spoken L2 (second language) learner English, collected from speaking tests performed on the Speak & Improve (S&I) learning platform [1]. The Speak & Improve Corpus 2025 consists of around 315 hours of recordings of L2 English learner speech on open speaking tasks, annotated with English speaking proficiency scores. In addition a subset of the data, around 55 hours, was manually transcribed, including disfluencies, and grammatically error corrected phrases [2]. The S&I platform was used by speakers from around the globe yielding a wide range of L1 (first language) backgrounds. The proficiency levels of the speakers range from Elementary (A2) to Advanced (C1) on the CEFR scale [3]. This breadth of speaker attributes and abilities is crucial for the development of inclusive language learning technologies. We believe this is the most comprehensive L2 English learner public corpus released to date.
@@ 4.2 Data distribution and selection
The Dev and Eval test sets were first defined to consist of full test submissions only i.e. where the learner has completed the test and all parts were successfully annotated at each phase of interest. 300 fully annotated test submissions were selected for each of the Dev and Eval test sets. These were extended for spoken language assessment (SLA) with a further (approximately) 150 submissions which had undergone phase 1 holistic proficiency scoring only. The submissions were selected so that the grade distribution was as even as possible but also reflects the nature of the learners using the platform. Table 4 shows the grade distribution for the Dev set. As can be seen the majority are in the CEFR range B1-B2+. The Eval and Train sets have similar grade distributions. The scores in Table 4 correspond to the marks awarded for a CEFR grade level on a part part basis. The overall score is the average of the scores across all the test parts so has a finer gradation. For this corpus the test is considered to consist of 4 parts. There was insufficient L1 and speaker data provided to incorporate this information into the data set selection. Note, this does mean that a speaker could appear in both the training and test sets. Table 5 shows the statistics for the Dev and Eval sets.
@@
Submissions Utterances Trans GEC SLA Transcript   GEC
Dev 438 5616 22.9 20.8 35.3 140k 105k
Eval 442 5642 22.7 20.4 35.4 140k 104k
@@
Submissions Utterances Trans GEC SLA Transcript   GEC
1 3068 18072 7.3 3.9 70.3 47k 22k
3 3060 3060 5.8 2.9 47.0 37k 14k
4 3005 3005 5.8 1.5 45.5 36k 7k
5 3085 15353 9.4 4.8 81.4 60k 24k
Total 6640 39490 28.2 13.0 244.2 170k 65k
@@ 5 Conclusions and Future Work
The S&I Corpus 2025 consists of audio and manual annotations of open speaking test submissions performed on the Speak & Improve learning platform. We provide around 315 hours of speech data from L2 English speakers at A2 to C1 CEFR levels, split into train, dev and eval data sets. There are around 950 fully annotated test submissions and the equivalent of a further 2500+ test submissions with manual speaking proficiency assessment scores.